OpenStudy (anonymous):

An elevator is moving downward with an acceleration of 2.0 m/s2. What is the force exerted on the elevator floor by a 75 kg person.

OpenStudy (anonymous):

\[R=m(g-a)\]

OpenStudy (anonymous):

\[= 75 (10-2) = 75*8 = 600N\] i took g as 10 if u take it as 9.8 u will get value near to this

OpenStudy (anonymous):

Thank you ! (:

OpenStudy (anonymous):

Welcome...if u r satisfied Give medals by clicking best answer option